OpenStudy (anonymous):

does anyone know the area between the curves y=(x^2)-x and y=x^3-4x^2+3x

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

like whats the area when refering to calculus

OpenStudy (anonymous):

the area will be the integral from 0 to 1 of (x^3-4x^2+3x)-(x^2-x) dx

OpenStudy (anonymous):

that's actually a really fun integral so i'm gonna compute it for you. lol. \[\int\limits_{0}^{1}(x^3-4x^2+3x)-(x^2-x)dx \] \[= x^4/4-4x^3/3+3x^2/2-x^3/3+x^2/2\] now evaluate from 0 to 1

OpenStudy (zarkon):

there is one more integral...they cross twice http://alturl.com/kg6eh

OpenStudy (anonymous):

thanks